OpenStudy (anonymous):

1.5 cos(5t)-0.8 sin(4t)=0 hey guys i got no idea how to solve this, can anyone please help me ? :)

OpenStudy (vishweshshrimali5):

\[\large{1.5\cos(5t) - 0.8\sin(4t) = 0}\] Multiply both sides by 10, we get: \[\large{15\cos(5t) - 8\sin(4t) = 0}\] Any doubt ?

OpenStudy (vishweshshrimali5):

Now, lets find out this: \[\large{\sqrt{15^2 + 8^2}}\] Can you tell me, what is the value of the above expression ?

OpenStudy (vishweshshrimali5):

@Wazzabob you there ?

OpenStudy (vishweshshrimali5):

Okay... I would try to be as detailed as possible and would leave some things for you to solve: I assume you have figured out \(\large{\sqrt{15^2 + 8^2}}\) which let is `a` (You must replace it with whatever value you got)

OpenStudy (vishweshshrimali5):

Now, divide both LHS and RHS by `a`: \[\large{15\cos(5t) - 8\sin(4t) = 0}\] \[\large{\implies {\cfrac{15}{a} \cos(5t) - \cfrac{8}{a} \sin(4t) = 0}}\] Any doubt till this step ?

OpenStudy (vishweshshrimali5):

Just to tell you, ` a = 17`

OpenStudy (vishweshshrimali5):

Okay lets continue... Now, this is an important and tricky step, so pay very close attention.... `If for some set of numbers 'a' and 'b'`, \(\large{a^2 + b^2 = 1}\), `then, I can replace 'a' and 'b' with : sin(u) and cos(u)`. Here: \[\large{u = \sin^{-1}(a) = \cos^{-1}(b)}\] Here also, \[\large{(\cfrac{15}{17})^2 + (\cfrac{8}{17})^2 = 1}\] So, let, \[\large{\cfrac{15}{17} = \sin u}\] and \[\large{\cfrac{8}{17} = \cos u}\] Thus, our initial equation becomes: \[\large{\sin u *\cos(5t) - \cos u *\sin (4t) = 0}\]

OpenStudy (vishweshshrimali5):

Now recall this equation: \[\large{\sin x \cos y = \cfrac{1}{2}(\sin(x+y) + \sin(x-y))}\] So: \[\large{\sin u *\cos(5t) = \cfrac{1}{2}(\sin(u+5t) + \sin(u-5t))}\] Similarly: \[\large{\sin(4t) * \cos(u) = \cfrac{1}{2}(\sin(4t+u) + \sin(4t-u))}\]

OpenStudy (vishweshshrimali5):

Thus, our equation becomes: \[\large{\cfrac{1}{2}[\sin(u+5t) + \sin(u-5t)] = \cfrac{1}{2}[\sin(4t+u) + \sin(4t-u)]}\] \[\large{\sin(u+5t) - \sin(u + 4t) = \sin(4t-u) - \sin(u-5t)}\] Now there is an another identity: \[\large{\sin x - \sin y = 2 \sin((x-y)/2) \cos((x+y)/2)}\] \[\large{\sin(u+5t) - \sin(u+4t) = 2\sin(t/2) \cos(u + 9t/2)}\] \[\large{\sin(4t-u) - \sin(u-5t) = 2\sin(9t/2 - u) \cos(-t/2)}\] Remember that: \[\large{\cos(-x) = \cos(x)}\] So, our equation finally becomes: \[\large{\sin(t/2) \cos(u + 9t/2) = \sin(9t/2 - u) \cos(t/2)}\]

OpenStudy (vishweshshrimali5):

I am terribly stuck :( I don't think, I can help you out after this... But, can you please recheck your question ?

OpenStudy (vishweshshrimali5):

One way to solve this is to write sin(5t) and cos(4t) in terms of sin(t) and cos(t) only and then solving the equation using a calculator...

OpenStudy (vishweshshrimali5):

Using this method, you would get: \[\cfrac{15}{17}\cos^5 (t) - \cfrac{150}{17} \sin^{2} (t) \cos^3(t) + \\ \cfrac{75}{17} \sin^4 (t) \cos(t) = \cfrac{32}{17}\sin(t)\cos^3(t) - \cfrac{32}{17}\sin^3(t) \cos(t)\] If I replace \(\sin(t)\) and \(\cos(t)\) by `a` and `b`, then you will have a system of equations: \[15b^5 - 150 a^2 b^3 + 75 a^4 b = 32ab^3 - 32a^3 b\tag{1}\] \[a^2 + b^2 = 1\tag{2}\] You can use an advanced equation solver to solve this set of equations and get the required answer

OpenStudy (vishweshshrimali5):

As you can see... it is very very lengthy... but solvable... That's the best I can do (at the very moment)... Hope this helps

OpenStudy (turingtest):

wow what a headache. This problem looks so benign and has such a pretty answer, too...

OpenStudy (vishweshshrimali5):

Yep

OpenStudy (anonymous):

the question is good, im very bad at solving these stuff :(